# medDreamer: Model-Based Reinforcement Learning with Latent Imagination on Complex EHRs for Clinical Decision Support

### v2

@@ 1 Introduction
Critical decision making in healthcare often involves managing highly dynamic and uncertain patient trajectories. Conditions such as sepsis or acute respiratory failure require timely and personalized interventions. However, the high complexity and heterogeneity of physiological responses challenge standardized care protocols. RL offers a principled framework for learning adaptive treatment strategies from complex data as it mirrors the sequential clinical decision-making process and directly optimizes long-term patient outcomes. Unlike standard supervised learning algorithms that aim to mimic historical clinician actions, RL can, in theory, surpass suboptimal clinical practices by learning decision treatment policies that are better aligned with long-term outcome-driven objectivesBanumathi et al. (2025).
@@
However, despite RL’s promise and increased interest, deploying RL solutions in the real-world clinical setting remains limited. The unique characteristics of EHR data, such as irregular sampling and highly sparse observations, pose significant challenges for standard RL methods. Clinical events occur sporadically based on the patient state and clinician decisions, which leads to the fact EHR data are highly irregular and often showcase informative missingnessRubin (1976), where the absence of measurements conveys key clinical information about patient conditions and dynamics. Existing solutions commonly rely on coarse discretization of patient trajectories into regular time windows (e.g., 4 hourly), and erroneous imputations using heuristics such as mean or median Komorowski et al. (2018). This could obscure important underlying patient dynamics, introduce information loss and modeling bias.
@@
Additionally, unlike traditional RL methods applied in gaming or robotics Goldwaser and Thielscher (2020); Yang and Wang (2020), real-world exploration during policy training is ethically infeasible, as experimentation on real patients is not an option. Therefore, offline RL is utilized in most solutions where we completely learn from retrospective data. Model-free offline RL methods have become the de facto in clinical decision-making tasks due to their algorithmic simplicity and minimal assumptions about environment dynamics. However, these solutions are tightly coupled to the historical actions without the ability to simulate or explore alternative decision paths that can help generalize beyond the potentially suboptimal and biased past clinical actions Nambiar et al. (2023); Luo et al. (2024a); Amirahmadi et al. (2023). This severely hinders the widely used model-free offline RL solutions from capturing the complexity of clinical trajectories and recommending robust, outcome-improving treatments in critical healthcare environments. To overcome these challenges, a fundamentally different approach is required that can model patient dynamics more faithfully, allowing counterfactual reasoning and supporting policy learning beyond what clinicians have historically practiced, where applicable.
@@
Accordingly, we propose medDreamer, a novel model-based RL framework designed to address the dual challenges of irregular clinical data and the constraints of model-free offline policy learning. medDreamer consists of a learned patient world model and a two-phase hybrid policy training pipeline. The world model captures patient dynamics from irregular observations and time-elapsed intervals using a novel Adaptive Feature Integration (AFI) Module, and generates simulated patient trajectories via latent imaginationHafner et al. (2024); Burchi and Timofte (2025); Samsami et al. (2024). By modeling trajectories directly in the latent space, medDreamer avoids the inaccuracies introduced by conventional time discretization and error-prone imputations, allowing more accurate next state prediction in the sparse observational data. Importantly, medDreamer supports rich policy exploration entirely offline. By sampling trajectories from the patient world model, we enable rich policy exploration while respecting the ethical boundaries. To ensure policy safety and alignment, we introduce a two-phase training strategy. Unlike the common practice of model-based RL approaches that purely train on imagined data Hafner et al. (2024), policies are initially trained on hybrid trajectories that contain both real and imagined steps, and then fine-tuned on fully simulated trajectories. We evaluate our solution on two clinical decision-making tasks: sepsis treatment recommendation and mechanical ventilation (MV) management, both leading causes of ICU mortality where optimal treatment policies remain an open questionNauka et al. (2025); Goligher et al. (2016). These challenges call for an effective data-driven solution capable of generalizing beyond limited retrospective clinical data to learn better treatment policies. Our main contributions are as follows:
@@ 2 Related Work
Inspired by the success in domains such as robotics Tang et al. (2025), game playing Mnih et al. (2013); Szita (2012) and autonomous driving Peng et al. (2024), RL has been widely adopted for various healthcare treatment recommendation tasks Ali (2022) for both chronic Perera et al. (2025); Yu et al. (2021a) and acute Jia et al. (2020); Raghu et al. (2017) treatment recommendation tasks. In this section, we review the literature from two perspectives: representation learning for patient state modeling and the evolution of model-based RL for treatment recommendation.
@@
Representation learning for RL in treatment recommendation: Accurate patient state representation is crucial for applying RL in clinical decision support. Several works have explored differential equation-based approaches to effectively represent irregular patient dynamics. GRU-ODE-BayesDe Brouwer et al. (2019) models sporadic EHR data using ordinary differential equations (ODEs), and subsequent work extends this for modeling treatment effects over timeDe Brouwer et al. (2022). Others utilize controlled differential equations (CDEs)Killian et al. (2020); Kidger et al. (2020); Morrill et al. (2021); Seedat et al. (2022) to track how patient states evolve. However, these methods build a continuous path over irregular states and are not suitable for creating discrete time latent trajectories. These methods also focus on irregularities for a single feature across time but do not account for the correlation among different features. In contrast, our approach is able to analyze both inter-feature and intra-feature dynamics to construct robust patient representations from sparse EHR data for downstream latent dynamics modeling and policy optimization.
@@
Model-based RL for treatment recommendation. The majority of past solutions utilize model-free RL, where treatment policies are learned directly from retrospective EHRs without an explicit model of patient dynamics Liu et al. (2024); Kondrup et al. (2023); Estiri and Mirinejad (2024); Tu et al. (2025). While efficient given enough data, model-free RL often suffers from sample inefficiency and limited capacity to reason about counterfactual outcomes since interactions with the real environment are not possible under clinical settings. In contrast, model-based reinforcement learning (MBRL) offers a principled framework for optimizing treatment strategies by simulating patient dynamics. Many previous works in this area build transition models based on Bayesian Neural Networks (BNNs). Killian et al.Killian et al. (2017) and Yao et al.Yao et al. (2018) propose HiP-MDPs to adapt transition dynamics based on latent patient embeddings for personalized decision-making. Raghu et al.Raghu et al. (2018) utilize a BNN to model transition dynamics for sepsis treatment. Parbhoo et al.Parbhoo et al. (2017) apply Gaussian process kernels for HIV treatment optimization, and Peng et al.Peng et al. (2018) extend similar methods to sepsis management. However, these methods operate directly in the real observation space, making it difficult to learn under noisy and irregular EHR data. Instead, our work infers the next state in the latent space by incorporating latent imaginationHafner et al. (2019, 2020, 2024). Further distinguishing our approach, unlike traditional model-based approaches that rely solely on world model-generated trajectories for policy learningSchrittwieser et al. (2020); Levine et al. (2016), we employ a two-phased approach to first ground the policy in real trajectories before introducing simulated rollouts. Previous offline MBRL methods, such as COMBOYu et al. (2021b) and MOPOYu et al. (2020), also use both real and simulated data uniformly during training. However, these approaches lack explicit grounding and exploration phases, and operate directly in raw observation spaces. Our proposed approach, medDreamer, addresses these limitations by introducing a structured, two-phased policy learning approach. We rely more on historical data during the first phase and shift to full exploration during the second phase, ensuring clinical reliability for healthcare applications.
@@ 3.3 Reward Function
For sepsis treatment, we define the intermediate reward $R^{\text{Sep}}_{\text{im}}(s_{t},s_{t+1})$ using changes in the Sequential Organ Failure Assessment (SOFA) score and lactate levelsRaghu et al. (2017); Komorowski et al. (2018):
@@ 3.4 Adaptive Feature Integration (AFI)
Clinical time series are often characterized by irregular sampling and feature sparsity. To effectively address these issues without erroneous discretization and imputations, we design an AFI module which treats missingness as a structural signal. The AFI module builds on a Factorization Machine (FM)Rendle (2010) based architecture to explicitly model both intra-feature temporal dynamics and inter-feature interactions across variables. Specifically AFI computes higher-order interactions among these features that extract robust representations from partially observed EHR data while preserving clinically meaningful patterns that helps the downstream policy learning task.
To keep the information within missing values, we encode both the observed feature values and their temporal context into a unified input representationChe et al. (2018). Suppose that we have time series $O=(o_{1},o_{2},\ldots,o_{T})\in\mathbb{R}^{T\times D}$ of length $T$ and dimension $D$. At time step $t$, the mask vector is defined as,
@@ 3.5.1 Posterior and Prior Dynamics:
Both $q_{\phi}$ and $p_{\theta}$ are parametrized as GRU network based sequential models due to its ability to capture long-term dependencies. The training objective for the transition model (prior) (see Equation 10) and representation model (posterior) (see Equation 9) is to maximize the evidence lower bound (ELBO)Jordan et al. (1999). This is achieved by minimizing the Kullback-Leibler (KL) divergence between the prior and the posterior distribution as,
@@ 3.6.1 Phase 1, Clinically Grounded Policy Initialization via Hybrid States:
We optimize the actor using the REINFORCE algorithmWilliams (1992) with a learned baseline. The actor loss is defined as:
@@ 4.1.1 Datasets:
For sepsis treatment, we use the large-scale MIMIC-IV Johnson et al. (2020) database, extracting 21,233 adult septic patients. We exclude patients with missing mortality outcomes or demographics, those with treatment withdrawal or multiple ICU stays. Each trajectory starts 24 hours before diagnosis and ends at 48 hours post-diagnosis. Each observation has 40 physiological features, and the outcome is 90-day mortality. For MV management, we use the eICU Collaborative Research (eICU)Pollard et al. (2018) database, and extract patients aged $\geq$16 who received MV for at least 24 hours. We also exclude patients with missing mortality outcomes or demographics, those ventilated for $>$14 days, and those with missing interventions. After filtering, 21,595 patients were retained with 41 physiological features, and the outcome is hospital mortality. See Appendix sectionB for full lists of features and actions.
@@ 4.1.2 Baselines:
We evaluate medDreamer against a range of model-based and model-free RL baselines on both tasks. We take reference from current benchmarks for treatment recommendation Luo et al. (2024b, a); Hargrave et al. (2024) and included naive baseline Random for sanity check, and a supervised learning Clinician Policy. For model-free baselines, we selected the ones that frequently appear in the benchmarks including Double Deep Q-Network (DDQN) Raghu et al. (2017), EZ-Vent Liu et al. (2024) based on Batch-Constrained Q-Learning (BCQ) Fujimoto et al. (2019) and DeepVent Kondrup et al. (2023) based on Conservative Q-Learning (CQL) Kumar et al. (2020). For model-based baselines, we choose a representative world model via latent imagination DreamerV3 Hafner et al. (2024) which is the latest version of Dreamer frameworkHafner et al. (2019), MBRL-BNN Raghu et al. (2018) which uses a BNN to model patient dynamics for sepsis treatment and MBRL-GMM Li et al. (2022) based Gaussian Mixture Model (GMM) and creates a surrogate world model using retrospective data for glucose control in DKA patients.
@@
EZ-VentLiu et al. (2024)(BCQ) 10.13 0.02 0.07 18.47 [18.45, 18.49] 7.09 -0.10 0.07 13.78 [13.76, 13.8]
DeepVentKondrup et al. (2023)(CQL) 10.42 0.03 0.07 19.39 [19.38, 19.39] 7.71 -0.10 0.08 13.8 [13.79 13.8]
DDQNRaghu et al. (2017) 10.10 0.02 0.07 17.77 [17.75, 17.79] 7.85 -0.10 0.08 21.5 [21.43, 21.56]
MBRL-BNNRaghu et al. (2018) 4.90 0.00 0.01 16.87 [14.5, 18.53] 6.34 0.03 0.08 16.7 [13.77, 19.62]
MBRL-GMMLi et al. (2022) -0.74 -0.01 -0.01 17.53 [17.51, 17.55] 4.09 -0.05 0.04 16.8 [16.79, 16.81]
DreamerHafner et al. (2024) 6.85 0.20 0.13 15.37 [13.13, 17.61] 12.95 0.28 0.20 13.96 [13.91, 16.57]
@@ 4.1.3 Evaluation Metrics:
Since we are unable to collect feedback from real patients, we evaluate our policy using multiple off-policy evaluation (OPE) metrics based on importance sampling Tokdar and Kass (2010) that estimates expectations under trained policy using samples from the behavior policy, including WIS, WPDIS, CWPDIS and ESS Thomas (2015). We also plot the mortality against expected return Wang et al. (2018) when executing clinicians’ actions. This plot helps us to infer the estimated mortality by matching the average return of recommended actions. We calculate the difference between dosages recommended by RL policy and clinician policy and plot it against the mortality. A good policy should be a v-shape, indicating that the mortality is lower when clinicians agree with our RL agent and higher otherwise.
@@ 4.2.2 RQ2: How does the learned policy align/differ from clinician decisions in treatment strategies?
The SOFA score is widely used in sepsis care to assess organ dysfunction, where higher scores indicate increased mortality risk. We observe the policy behavior across varying clinical severity by categorizing the patient into 4 SOFA ranges: Low (5$\leq$ 5≤ 5), Mid (6-14), High (15$\geq$ 15≥ 15), and All. We compare sepsis action distributions across SOFA severity levels to assess treatment adaptation among baseline models (see Figure 4). Clinicians displayed clear progression from conservative to aggressive interventions with increasing severity. Our method generally follows this trend, for example, in its increased use of vasopressors under high SOFA conditions. However, medDreamer is highly conservative in the use of IV fluid, providing a new perspective on decreasing IV fluid dosage. In contrast, model-free baselines like DDQN and EZ-Vent exhibit more diffuse and less interpretable action patterns, suggesting potential instability or lack of calibration. Other model-based methods, such as MBRL-GMM and MBRL-BNN, favor consistent low-level actions but lack sufficient flexibility to adjust to patient severity. These results demonstrate medDreamer’s ability to learn severity-aware treatment strategies that align more closely with clinical patterns.

### v3

@@ 1 Introduction
Critical decision making in healthcare often involves managing highly dynamic and uncertain patient trajectories. Conditions such as sepsis or acute respiratory failure require timely and personalized interventions. However, the high complexity and heterogeneity of physiological responses challenge standardized care protocols. RL offers a principled framework for learning adaptive treatment strategies from complex data as it mirrors the sequential clinical decision-making process and directly optimizes long-term patient outcomes. Unlike standard supervised learning algorithms that aim to mimic historical clinician actions, RL can, in theory, surpass suboptimal clinical practices by learning decision treatment policies that are better aligned with long-term outcome-driven objectives[banumathi2025reinforcement].
@@
However, despite RL’s promise and increased interest, deploying RL solutions in the real-world clinical setting remains limited. The unique characteristics of EHR data, such as irregular sampling and highly sparse observations, pose significant challenges for standard RL methods. Clinical events occur sporadically based on the patient state and clinician decisions, which leads to the fact EHR data are highly irregular and often showcase informative missingness[rubin1976inference], where the absence of measurements conveys key clinical information about patient conditions and dynamics. Existing solutions commonly rely on coarse discretization of patient trajectories into regular time windows (e.g., 4 hourly), and erroneous imputations using heuristics such as mean or median [komorowski2018artificial]. This could obscure important underlying patient dynamics, introduce information loss and modeling bias.
@@
Additionally, unlike traditional RL methods applied in gaming or robotics [goldwaser2020deep, yang2020overview], real-world exploration during policy training is ethically infeasible, as experimentation on real patients is not an option. Therefore, offline RL is utilized in most solutions where we completely learn from retrospective data. Model-free offline RL methods have become the de facto in clinical decision-making tasks due to their algorithmic simplicity and minimal assumptions about environment dynamics. However, these solutions are tightly coupled to the historical actions without the ability to simulate or explore alternative decision paths that can help generalize beyond the potentially suboptimal and biased past clinical actions [nambiar2023deep, luo2024reinforcement, amirahmadi2023deep]. This severely hinders the widely used model-free offline RL solutions from capturing the complexity of clinical trajectories and recommending robust, outcome-improving treatments in critical healthcare environments. To overcome these challenges, a fundamentally different approach is required that can model patient dynamics more faithfully, allowing counterfactual reasoning and supporting policy learning beyond what clinicians have historically practiced, where applicable.
@@
Accordingly, we propose medDreamer, a novel model-based RL framework designed to address the dual challenges of irregular clinical data and the constraints of model-free offline policy learning. medDreamer consists of a learned patient world model and a two-phase hybrid policy training pipeline. The world model captures patient dynamics from irregular observations and time-elapsed intervals using a novel Adaptive Feature Integration (AFI) Module, and generates simulated patient trajectories via latent imagination[hafner2024masteringdiversedomainsworld, burchi2025learning, samsami2024mastering]. By modeling trajectories directly in the latent space, medDreamer avoids the inaccuracies introduced by conventional time discretization and error-prone imputations, allowing more accurate next state prediction in the sparse observational data. Importantly, medDreamer supports rich policy exploration entirely offline. By sampling trajectories from the patient world model, we enable rich policy exploration while respecting the ethical boundaries. To ensure policy safety and alignment, we introduce a two-phase training strategy. Unlike the common practice of model-based RL approaches that purely train on imagined data [hafner2024masteringdiversedomainsworld], policies are initially trained on hybrid trajectories that contain both real and imagined steps, and then fine-tuned on fully simulated trajectories. We evaluate our solution on two clinical decision-making tasks: sepsis treatment recommendation and mechanical ventilation (MV) management, both leading causes of ICU mortality where optimal treatment policies remain an open question[nauka2025challenges, goligher2016clinical]. These challenges call for an effective data-driven solution capable of generalizing beyond limited retrospective clinical data to learn better treatment policies. Our main contributions are as follows:
@@ 2 Related Work
Inspired by the success in domains such as robotics [tang2025deep], game playing [mnih2013atari, Szita2012] and autonomous driving [peng2024improving], RL has been widely adopted for various healthcare treatment recommendation tasks [ali2022reinforcement] for both chronic [si_25, yu2021reinforcement] and acute [jia2020safe, raghu2017deep] treatment recommendation tasks. In this section, we review the literature from two perspectives: representation learning for patient state modeling and the evolution of model-based RL for treatment recommendation.
@@
Representation learning for RL in treatment recommendation: Accurate patient state representation is crucial for applying RL in clinical decision support. Several works have explored differential equation-based approaches to effectively represent irregular patient dynamics. GRU-ODE-Bayes[de2019gru] models sporadic EHR data using ordinary differential equations (ODEs), and subsequent work extends this for modeling treatment effects over time[de2022predicting]. Others utilize controlled differential equations (CDEs)[killian2020empirical, kidger2020neural, morrill2021neural, seedat2022continuous] to track how patient states evolve. However, these methods build a continuous path over irregular states and are not suitable for creating discrete time latent trajectories. These methods also focus on irregularities for a single feature across time but do not account for the correlation among different features. In contrast, our approach is able to analyze both inter-feature and intra-feature dynamics to construct robust patient representations from sparse EHR data for downstream latent dynamics modeling and policy optimization.
@@
Model-based RL for treatment recommendation. The majority of past solutions utilize model-free RL, where treatment policies are learned directly from retrospective EHRs without an explicit model of patient dynamics [liu2024reinforcement, kondrup2023towards, estiri2024model, tu2025offline]. While efficient given enough data, model-free RL often suffers from sample inefficiency and limited capacity to reason about counterfactual outcomes since interactions with the real environment are not possible under clinical settings. In contrast, model-based reinforcement learning (MBRL) offers a principled framework for optimizing treatment strategies by simulating patient dynamics. Many previous works in this area build transition models based on Bayesian Neural Networks (BNNs). Killian et al.[killian2017robust] and Yao et al.[yao2018direct] propose HiP-MDPs to adapt transition dynamics based on latent patient embeddings for personalized decision-making. Raghu et al.[raghu2018model] utilize a BNN to model transition dynamics for sepsis treatment. Parbhoo et al.[parbhoo2017combining] apply Gaussian process kernels for HIV treatment optimization, and Peng et al.[peng2018improving] extend similar methods to sepsis management. However, these methods operate directly in the real observation space, making it difficult to learn under noisy and irregular EHR data. Instead, our work infers the next state in the latent space by incorporating latent imagination[hafner2019dream, hafner2020mastering, hafner2024masteringdiversedomainsworld]. Further distinguishing our approach, unlike traditional model-based approaches that rely solely on world model-generated trajectories for policy learning[schrittwieser2020mastering, levine2016end], we employ a two-phased approach to first ground the policy in real trajectories before introducing simulated rollouts. Previous offline MBRL methods, such as COMBO[yu2021combo] and MOPO[yu2020mopo], also use both real and simulated data uniformly during training. However, these approaches lack explicit grounding and exploration phases, and operate directly in raw observation spaces. Our proposed approach, medDreamer, addresses these limitations by introducing a structured, two-phased policy learning approach. We rely more on historical data during the first phase and shift to full exploration during the second phase, ensuring clinical reliability for healthcare applications.
@@ 3.3 Reward Function
For sepsis treatment, we define the intermediate reward $R^{\text{Sep}}_{\text{im}}(s_{t},s_{t+1})$ using changes in the Sequential Organ Failure Assessment (SOFA) score and lactate levels[raghu2017deep, komorowski2018artificial]:
@@ 3.4 Adaptive Feature Integration (AFI)
Clinical time series are often characterized by irregular sampling and feature sparsity. To effectively address these issues without erroneous discretization and imputations, we design an AFI module which treats missingness as a structural signal. The AFI module builds on a Factorization Machine (FM)[rendle2010factorization] based architecture to explicitly model both intra-feature temporal dynamics and inter-feature interactions across variables. Specifically AFI computes higher-order interactions among these features that extract robust representations from partially observed EHR data while preserving clinically meaningful patterns that helps the downstream policy learning task.
To keep the information within missing values, we encode both the observed feature values and their temporal context into a unified input representation[che2018recurrent]. Suppose that we have time series $O=(o_{1},o_{2},\ldots,o_{T})\in\mathbb{R}^{T\times D}$ of length $T$ and dimension $D$. At time step $t$, the mask vector is defined as,
@@ 3.5.1 Posterior and Prior Dynamics:
Both $q_{\phi}$ and $p_{\theta}$ are parametrized as GRU network based sequential models due to its ability to capture long-term dependencies. The training objective for the transition model (prior) (see Equation 10) and representation model (posterior) (see Equation 9) is to maximize the evidence lower bound (ELBO)[jordan1999introduction]. This is achieved by minimizing the Kullback-Leibler (KL) divergence between the prior and the posterior distribution as,
@@ 3.6.1 Phase 1, Clinically Grounded Policy Initialization via Hybrid States:
We optimize the actor using the REINFORCE algorithm[williams1992simple] with a learned baseline. The actor loss is defined as:
@@ 4.1.1 Datasets:
For sepsis treatment, we use the large-scale MIMIC-IV [johnson2020mimic] database, extracting 21,233 adult septic patients. We exclude patients with missing mortality outcomes or demographics, those with treatment withdrawal or multiple ICU stays. Each trajectory starts 24 hours before diagnosis and ends at 48 hours post-diagnosis. Each observation has 40 physiological features, and the outcome is 90-day mortality. For MV management, we use the eICU Collaborative Research (eICU)[pollard2018eicu] database, and extract patients aged $\geq$16 who received MV for at least 24 hours. We also exclude patients with missing mortality outcomes or demographics, those ventilated for >>14 days, and those with missing interventions. After filtering, 21,595 patients were retained with 41 physiological features, and the outcome is hospital mortality. See Appendix sectionB for full lists of features and actions.
@@ 4.1.2 Baselines:
We evaluate medDreamer against a range of model-based and model-free RL baselines on both tasks. We take reference from current benchmarks for treatment recommendation [luo2024dtr, luo2024reinforcement, hargrave2024epicare] and included naive baseline Random for sanity check, and a supervised learning Clinician Policy. For model-free baselines, we selected the ones that frequently appear in the benchmarks including Double Deep Q-Network (DDQN) [raghu2017deep], EZ-Vent [liu2024reinforcement] based on Batch-Constrained Q-Learning (BCQ) [fujimoto2019off] and DeepVent [kondrup2023towards] based on Conservative Q-Learning (CQL) [kumar2020conservative]. For model-based baselines, we choose a representative world model via latent imagination DreamerV3 [hafner2024masteringdiversedomainsworld] which is the latest version of Dreamer framework[hafner2019dream], MBRL-BNN [raghu2018model] which uses a BNN to model patient dynamics for sepsis treatment and MBRL-GMM [li2022electronic] based Gaussian Mixture Model (GMM) and creates a surrogate world model using retrospective data for glucose control in DKA patients.
@@
EZ-Vent[liu2024reinforcement](BCQ) 10.13 0.02 0.07 18.47 [18.45, 18.49] 7.09 -0.10 0.07 13.78 [13.76, 13.8]
DeepVent[kondrup2023towards](CQL) 10.42 0.03 0.07 19.39 [19.38, 19.39] 7.71 -0.10 0.08 13.8 [13.79 13.8]
DDQN[raghu2017deep] 10.10 0.02 0.07 17.77 [17.75, 17.79] 7.85 -0.10 0.08 21.5 [21.43, 21.56]
MBRL-BNN[raghu2018model] 4.90 0.00 0.01 16.87 [14.5, 18.53] 6.34 0.03 0.08 16.7 [13.77, 19.62]
MBRL-GMM[li2022electronic] -0.74 -0.01 -0.01 17.53 [17.51, 17.55] 4.09 -0.05 0.04 16.8 [16.79, 16.81]
Dreamer[hafner2024masteringdiversedomainsworld] 6.85 0.20 0.13 15.37 [13.13, 17.61] 12.95 0.28 0.20 13.96 [13.91, 16.57]
@@ 4.1.3 Evaluation Metrics:
Since we are unable to collect feedback from real patients, we evaluate our policy using multiple off-policy evaluation (OPE) metrics based on importance sampling [tokdar2010importance] that estimates expectations under trained policy using samples from the behavior policy, including WIS, WPDIS, CWPDIS and ESS [thomas2015safe]. We also plot the mortality against expected return [wang2018supervised] when executing clinicians’ actions. This plot helps us to infer the estimated mortality by matching the average return of recommended actions. We calculate the difference between dosages recommended by RL policy and clinician policy and plot it against the mortality. A good policy should be a v-shape, indicating that the mortality is lower when clinicians agree with our RL agent and higher otherwise.
@@ 4.2.2 RQ2: How does the learned policy align/differ from clinician decisions in treatment strategies?
The SOFA score is widely used in sepsis care to assess organ dysfunction, where higher scores indicate increased mortality risk. We observe the policy behavior across varying clinical severity by categorizing the patient into 4 SOFA ranges: Low ($\leq 5$), Mid (6-14), High ($\geq 15$), and All. We compare sepsis action distributions across SOFA severity levels to assess treatment adaptation among baseline models (see Figure 4). Clinicians displayed clear progression from conservative to aggressive interventions with increasing severity. Our method generally follows this trend, for example, in its increased use of vasopressors under high SOFA conditions. However, medDreamer is highly conservative in the use of IV fluid, providing a new perspective on decreasing IV fluid dosage. In contrast, model-free baselines like DDQN and EZ-Vent exhibit more diffuse and less interpretable action patterns, suggesting potential instability or lack of calibration. Other model-based methods, such as MBRL-GMM and MBRL-BNN, favor consistent low-level actions but lack sufficient flexibility to adjust to patient severity. These results demonstrate medDreamer’s ability to learn severity-aware treatment strategies that align more closely with clinical patterns.